OpenStudy (anonymous):

Find the first 5 terms of the sequence: a1 = 500, an = (an-1)/5.

OpenStudy (cornzyblack):

a1 = 500 a2 = ((500 * 2) - 1)/5 = 999/5 a3 = ((500 * 3) - 1)/5 =1499/5 a4 = ((500 * 4) - 1)/5 = 1999/5 a5 = ((500 * 5)- 1)/5 = 2499/5 Are You sure you typed the write thing that is an = (an - a)/5

OpenStudy (anonymous):

a_n

OpenStudy (cornzyblack):

i mean is the formula correct or is it (a(n-1))/5 ??

OpenStudy (anonymous):

It's right. just copied and pasted it

OpenStudy (cornzyblack):

then thats tha answer in ma first Comment

OpenStudy (anonymous):

Thank you